OpenStudy (juan1857):

Find each principal root. Express the result in the form a+bi with a and b rounded to the nearest hundredth. (32(cos2pi/3 +isin2pi/3))^1/5

OpenStudy (loser66):

I saw you post it twice, what don't you understand?

OpenStudy (juan1857):

oh I don't understand the last part of the answer

OpenStudy (loser66):

What is De Moivre's theorem?

OpenStudy (juan1857):

well I got up to 2(cos2pi/15 +isin2pi/15)

OpenStudy (loser66):

\(if, z=r*(a *cos x + i*bsinx)\) then \(z^{\color{red}{n}}= r^{\color{red}{n}}*(a cos(\color{red}{n}*x+i*b*sin(\color{red}{n}*x))\) Your n is 1/5, r = 32 a=1 b=1 x=2pi/3, just plug them all in \(z^{\color{red}{1/5}}= 32^{\color{red}{1/5}}* cos(\color{red}{1/5}*(2pi/3)+i*sin(\color{red}{1/5}*(2pi/3))\)

OpenStudy (loser66):

\(32^{\color{red}{1/5}}= \sqrt[\color{red}{5}]{32}=2\)

OpenStudy (loser66):

\(\color{red}{1/5}*2pi/3= 2pi/15\)